In a pair-matched case-control study of risk factors for carcinoma, dietary history of 30 patients with the disease and 30 control subjects were collected. Below is a table that summarizes their history of corn and corn oil consumption.

Pairs Cases - Carcinoma patients Controls - healthy subjects
1 Yes Yes
2 Yes Yes
3 Yes Yes
4 Yes Yes
5 Yes Yes
6 Yes Yes
7 Yes No
8 Yes No
9 Yes No
10 Yes No
11 Yes No
12 Yes No
13 Yes No
14 Yes No
15 Yes No
16 Yes No
17 Yes No
18 Yes No
19 Yes No
20 No No
21 No No
22 No No
23 No No
24 No No
25 No No
26 No Yes
27 No Yes
28 No Yes
29 No Yes
30 No Yes

a. What is estimated odds ratio? what is the 95% confidence interval of population odds ratio?

b. Perform hypothesis test and see if there is any difference in efficacy between two treatments. level of sig - 0.05. state null and alternate hypothesis, test statistic, p value and your conclusion.
Answer #1

This is the question of case-control data where the counts in this table(table structured using the information above) represent the number of pairs and not the number of individuals.

Cases/control Yes No
Yes 6 13
No 5 6

No. of Yes-yes pairs are 6, No. of Yes-No pairs are 13, No. of No-Yes pairs are 5, No. of No-No pairs are 6 (from the information provided in the question)

a) Here Yes-Yes and No-No are known as the concordant pairs of the sample and Yes-No and No-Yes are known as discordant pairs. Although there are N = 30 pairs but we are only interested in the discordant pairs which are u = 13 and v = 5.

Odds Ratio =

The confidence interval for odds-ratio =

where z = 1.96 at 95% level of confidence.

Now first let us find

=

= 0.526

Now we will calculate the CI for odds ratio=

=

= -0.076,1.99

Hence the 95% confidence interval for case control data's odd ratio is [-0.076,1.99].

b) Also, When the number of discordant pairs (u + v) is 10 or greater, you can test H0: OR = 1 with McNemar’s chi-square statistic.

Since u+v = 13+5 = 18

H0: OR = 1

H1: OR > 1

The regular and continuity-correct McNemar’s chi squares are shown below:

Because of the relation between the chi-square distributions and z distributions, the above formulas can be re-expressed: